OpenStudy (anonymous):

The area of a rectangular field is 160m^2 and its length is twice its width .Find the dimensions of the feild.

OpenStudy (anonymous):

A=LxW L=2W 160m^2= W x 2W 160m^2= 2W^2 160m^2/2 = 2W^2/2 80m^2=W^2 W=square root of 80 W=8.944m L=2W L=2(8.944) L=17.89m CHECKING: A=LxW 160m^2=8.944m x 17.89m 160m^2=159.99 or 160m^2

OpenStudy (anonymous):

\[A= 160m ^{2}\] A= L x W its length is twice its width L = 2w so write the equation... A = L x W 160m^2 = (2W) W 160m^2/2 = 2W^2/2 80m^2 = W^2 8.944 = W Solve for L or length L = 2W L = 2(8.944) L = 17.89m Check 160m^2=8.944m x 17.89m 160m^2=159.99 or 160m^2